Wall assembly 3 has the following components: Inside air film 0.5 inch gypsum board 5.5 inch insulation (?) 0.5 inch OSB (Oriented Strand Board) 1 inch stucco Outside air film The building code requires a maximum U-Value of 0.0436 for the wall assembly. What is the minimum R-Value required for the insulation to meet code?
Answer #1

Ans) Given,

Required U value = 0.0436 BTU / F ft2 hr

We know, R value = 1 / U-value

=> R-value = 1 / 0.0436 = 22.935 F ft2 hr / BTU

For 0.50 inch interior air film , R-value (R1) = 0.68

For 0.5 in gypsum board insulation, R value (R2) = 0.45

For 5.5 inch insulation  , R value (R3) = x

For 0.50 in OSB , R value (R4) = 0.62

For 1" outside air film , R value (R5) = 0.17

Total R- value = R1 + R2 + R3 + R4 + R5

=> 22.935 = 0.68 + 0.45 + x + 0.62 + 0.17

=> x = 21.015 F ft2 hr/ BTU

Hence, minimum required R value for insulation to meet code is 21.015 F ft2 hr/BTU